OpenStudy (dominirican1013):

Which of the following equations are identities? Check all that apply.

OpenStudy (dominirican1013):

\[A. \cot^2x+\csc^2x=1\] \[B. \tan^2x=\sec^2x-1\] \[C. \sin^2x=1-\cos^2x\] \[D. \sin^2x-\cos^2x=1\]

OpenStudy (dominirican1013):

I know for sure it is not D.

OpenStudy (dominirican1013):

@welshfella

OpenStudy (mertsj):

Do you know the Pythagorean Identities?

OpenStudy (dominirican1013):

I do the ones that I think are right is B and D but I am not too sure

OpenStudy (mertsj):

B & C

OpenStudy (dominirican1013):

oh okay thank you. so D is not correct

OpenStudy (mertsj):

\[\sin ^2+\cos ^2=1\]

OpenStudy (mertsj):

no

OpenStudy (dominirican1013):

aorry I did not mean to say D lol

OpenStudy (dominirican1013):

Thank you